Tech Staff Meeting Notes October 22, 2007
Current Top Priorities are: Station Maintenance. Mud-Island. Two new GPS stations. (Also wish to get INV populated.)
There are four levels of priority.
|Priorities||Wish to try to get these done (or make progress on) before the next meeting (yeah right)|
|Second Tier||Fit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month|
|As Time Permits||Fit these in when you can but not instead of first or second priorities|
|Other||Stuff we hope to get to and don't want to forget about|
The task list was updated July 7, 2007 to incorporate some tasks previously performed by Dave Greganti. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.
Station Maintenance, NMSZ
Station maintenance will become the top priority after the GPS work is finished around November 2.
Holly's list on 10/21/2007:
Isis Oxidation Problem
Need a dual mux assembly or two. Have the AMUX64T's. Pat will build up
the rest of the hardware.
MKTA spare air conditioner is repaired. Needs to be returned to the doghouse.
Need a remote reset at lnxt. Have permission from WASL to piggy back on theirs.
Just need Pat to wire it in and show us how to use it.
Need new RF preamps at MKTA.
MKTA also needs a GX620, ISIS telemetry upgrades, radio realign, etc.
PEBM and HENM sometimes have telemetry fades. Should reevaluate comms.
NMAD needs brush control when the leaves fall.
Mud Island vertical array is operational.
The number 3 recorder keeps losing gps lock.
The recorder disks need to be reformatted. This can't be done remotely. The constant read/write to the fat file systems is going to make this reformatting something that needs to be done periodically. Requires physically going under the bridge to do it. Dont ya just love dos and obsolete hardware?
The wireless resets seem to be getting noisy making it more tricky than normal to reset.
Lost contact with two of seven recorders.
We're in the process of getting James out of the station maintenance business. GPS stations will be treated as any of the seismic network stations. Mac will be the lead for routine maintenance. Data handling and processing is still done by Bob et al.
Greg is in charge of getting anything done to the GPS stations to Steinerize them. This is necessary to make sure we can maintain them without burning alot of time on reapplying bandaids. Note that they are in pretty good shape already. The primary hassle is communications.
Most of the work for the upgrades and two new stations will be complete in the next to weeks. Driller is currently scheduled to be on site the week of Halloween. Chris will need to be on-site. According to the bid, we do the install and the driller drills and assists. Bob should probably be there too. Chris is looking into how best to weld the stainless rods.
Comms for MIAR, CJTR.
Install 2 Monduments, LCHS and New Harkham.
Install LCHS electronics and fiber run (not fiber anymore thought there will be optical isolation).
Install New Harkham tower and instrument enclosure.
Install New Madrid end of link (4' dish).
Install RLAP tower and monument equipment.
Van Hill gate got bull dozed. Need to replace it when they're done
TZTN needs a gate.
NEIC requested John's help in siting the Campobello station (northeast South Carolina). He found a state park willing to host it. Need to collect some noise.
NEIC decided to fix Murphy.
BRBC needs a new XLAK GPS antenna. John has the antenna.
The map shows stations with yellow triangles being new ones.
|SMNC||4||done (see note above)|
TVNC is installed. The radios need repaired. John sent them to Greg.
Chris will schedule a trip with John the week after Halloween. Steve and Al will get 5 GX620's ready to go. These will go at Sewanee, two at Athens, Star Mountain, and Mt. Mitchell.
Routine Maintenance Kudzu, battery check, etc, etc.
Second Tier Priorities
CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready.
Steve spoke with Dean Kemp at NWS. They are willing to let us piggyback on their cable internet (not .nws.noaa.gov). Dean is looking into the possibility of a static IP. This would get NAIT online.
Lenox Agreement. Mitch spoke with Ron Martens, EMF Broadcasting. There is a desire for some sort of MOU and a visit to CERI. Mitch will set this up.
There is a new FDL Policy.
Bill reports that his laptop works well with the equipment but the others do not. Steve will show him how to get the others set up like his.
Steve has boot floppies and instructions for downloading data from the multi-channel seismograph without using the NT domain (which is defunct). The images and instructions are also available at /gaia/data/fdl/strataview.boot.floppy.
Bill is fixing the several breakout boxes and cables for the 6TD's that got damaged. Mitch will instruct people who check these out on the how to use the handle for a handle instead of the cable.
As Time Permits
3876 Power The power for the new lighting on the east side of the general parking lot to the west of CERI is coming from 3876. When the contractor is finished, Chris will need to check it to see if there is a way to isolate the circuits for the house from the circuits for the sidewalk lighting. Otherwise, Jim pointed out that it increases the risk of lighting surges that could damage equipment.
Chris is going to make a diagram for 3890. Good luck son.
Jim has all the broadband sensors defined and is working on additional
equipment (e.g. isis) and the connections. Mitch will install Bill's
account in mysql and inv so he can start building stations. Mitch needs
to sit down with Jim and figure out what we need to have ISTI help
us with. Sid will be available after about mid-May to finish up anything
we need his help to take care of. Apparently the downloaded docs from
guralp for the cmg5td's have the old +-4g contants. Mitch has paper
documentation with the correct +-2g constants.
This is languishing. It needs to become a priority for both Mitch and Jim in December.
Steiner Brain Dump Seminar Need more of them as weather permits. Especially calibration and computer room power.
Thin-wire aironets are no longer available. This will be a problem for the DBB. It will need to be addressed hopefully later rather than sooner. Steve reports the new ones play with the old ones so its a resolvable problem.
Withers needs to determine current orientation for MPHB so we can pull it and orient correctly.
There are a plethora of preventive maintenance type things around CERI that Greg handles. He will construct a list of them so we can decide who is doing what.
The infrasound array at the Langston farm was designed with a life of 6 months. It is not currently operational. Chuck would like to get it working again and possibly install others elsewhere. Probably need a redesign and rebuild for this to put back the corners that were cut for the temporary deployment.
Need to teach someone (our hazmat dude?) to neutralize etchant and clean the etching tanks. Along those same lines, Greg has a ton of drawings that play only with the "Eagle" software on his computer. Mitch is unaware of a way to convert to a more generic format. Also require a printer that does negative (the one on Greg's serial port does, others do not). Mitch knew this dependence on vendor specific software would bite us one day. That day is soon.
Pink Palace hardware needs to be rebuilt.
John needs a spectrum analyzer.
Evaluate computer room power. Steve will be the keeper of the spreadsheet post-Steiner.
Test worm rack batteries.
Install new netgear switches.
Build more battery regulators.